Archive for July, 2017

Incident: VoIP Disruption

Posted: Thursday, July 13th, 2017 at 10:46 by Richard Partridge

Engineers are currently investigating an incident that is causing a loss of service on our VoIP platform. Further information will be provided as and when it becomes available.

  1. Richard Partridge says:

    Service has now been restored. We apologise for any inconvenience.

Email Disruption

Posted: Thursday, July 13th, 2017 at 10:15 by Richard Partridge

We are currently receiving reports that a small number of DSL customers are experiencing difficulty when trying to send or receive emails. Our engineers are currently investigating and further information will be provided as and when it becomes available.

  1. Richard Partridge says:

    Engineers are still investigating the email issue at this time. Further updates will be provided over the course of the day.

  2. Richard Partridge says:

    Engineers are continuing their investigation. Further updates will be provided over the course of the day

  3. Richard Partridge says:

    Engineers have now been able to recreate the symptoms being reported and are investigating a possible root cause of the problem. Further updates will follow in due course.

  4. Richard Partridge says:

    Adjustments have been made to try and resolve this problem, however we will need to return to normal bandwidth levels in order to determine their effectiveness. Further updates will be provided tomorrow morning once the impact of these changes have been ascertained.

  5. Richard Partridge says:

    “Yesterday’s adjustments have had a positive affect, but haven’t entirely resolved the problem. We’re continuing to work on this and will provide further updates over the course of the day.”

  6. Richard Partridge says:

    No further issues have been reported over the course of the afternoon. We will closely monitor the situation over the course of the weekend and advise of any further occurrences of this issue accordingly.

BT circuit issue

Posted: Thursday, July 6th, 2017 at 10:56 by Adam Heath

We are aware of an issue affecting a number of BT Ethernet circuits which connect to our Columbo core.   Our initial diagnostics indicate the issue to be on BT’s hub and we have this raised with them already for investigation.

We will provide further information once it becomes available.

  1. Adam Heath says:

    We are continuing to correspond and troubleshoot with the suppliers. As a precaution, we have also tasked out an engineer to double check our equipment at the site.

    Further updates will be available shortly.

  2. Adam Heath says:

    We have restored service to affected circuits following an engineer replacing a faulty optic at site. We’re continuing to monitor from our side.